OpenStudy (anonymous):

2 1/4 - 1 5/8 Fractions

OpenStudy (anonymous):

-1/8 i think.... without the braces

OpenStudy (anonymous):

i got 3 3/8 Is this wrong?

OpenStudy (anonymous):

2*1/4=2/4=1/2 1*5/8=5/8 1/2-5/8=(8-10)/16=-2/16=-1/8 thats how i got there...!!

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

I thought they were mixed fractions. 2 1/4 - 1 5/8 = 9/4 - 13/8 = (18-13)/8 = 5/8

OpenStudy (anonymous):

@Aadarsh i didn't consider mixed... i took them as normal ones....

OpenStudy (anonymous):

@boomboomboom24 i guess you have to be a bit more specific

OpenStudy (anonymous):

Actually, its typing mistake. Should have used bracktes or some other symbols to demark properly.

OpenStudy (anonymous):

@Aadarsh thats why i mentioned i assumed them with braces in my first reply.. well i guess the both of us are right in our own worlds....!!

OpenStudy (anonymous):

Yeah sure Bhai.